A manufacturer of dog food is concerned about the low levels of sales recorded of one of its tinned


products and wants to consider removing this product from its line. Before doing so, a limited scale


market surveyed of 1000 customers that buy this particular brand of dog food is conducted. Of the


people surveyed, 230 said that they would like the manufacturer to keep the product under


investigation as one of the products in the manufacturer’s line. Use this information to construct the


95% confidence interval for the proportion of all customers of this manufacturing brand that would


like the manufacturer to keep the product under investigation as one of the products in the


manufacturer’s line.

1
Expert's answer
2021-11-01T17:37:11-0400

The sample proportion is computed as follows, based on the sample size "N = 1000" and the number of favorable cases "X = 230"


"\hat{p}=\dfrac{X}{N}=\dfrac{230}{1000}=0.23"

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}})"

"=(0.23-1.96\sqrt{\dfrac{0.23(1-0.23)}{1000}},"

"0.23+1.96\sqrt{\dfrac{0.23(1-0.23)}{1000}})"


"=(0.204,0.256)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.204 < p < 0.256," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.204, 0.256)."